OpenStudy (anonymous):

Which set of integers is a Pythagorean triple? A. 6, 8, 12 B. 5, 12, 13 C. 15, 9, 17 D. 3, 5, 6

OpenStudy (mathstudent55):

Do you know the Pythagorean theorem equation, \(a^2 + b^2 = c^2\) ?

OpenStudy (cwrw238):

try each one out eg 6^2 = 3^2 + 5^2 36 = 9 + 25 - not this one

OpenStudy (anonymous):

and thanks for not helping

OpenStudy (mathstudent55):

For each choice, replace c with the largest number, and "a" and b with the other two numbers. Whichever triple makes the equation true is it.